You have received a bonus of $5000 that you would like to invest. You have three options. BMO savings account which earns 15 percent compounded daily; td savings account which earns 15.5 percent compounded quarterly and cibc savings account which earns 16 percent compounded annually. Which would you chose and why???
Answer #1

Ans We choose td savings account which earns 15.5 percent compounded quarterly since it has highest Future Value.

1) BMO FV = PV (1 + r )^n
FV = 5000*(1+15%/365)^365
FV = 5808.99
2) TD FV = PV (1 + r )^n
FV = 5000*(1+15.5%/4)^4
FV = 5821.22
3) CIBIC FV = PV (1 + r )^n
FV = 5000*(1+16%)^1
FV = 5800.00